OpenStudy (anonymous):

A pair of shirt is 99$ dollar and take 10% off how much will the shirt will be

hero (hero):

99-9.9 = ?

OpenStudy (agentorange96):

99/100=0.99 (So 1% of 99 = 0.99) 0.99x10=9.9 (So 10% of 99 = 9.9) And here we are left with Hero's answer: 99-9.9=? Do the math...

hero (hero):

I started off by moving the decimal one place to the left and then subtracting that from 99

OpenStudy (agentorange96):

Oh, and if the question is how much is one shirt, just divide the answer by two, your wording is a bit confusing toward the end...